# Characterizing the Distinguishability of Product Distributions through Multicalibration

### v3

@@ 9.1 Acknowledgements
This project was inspired by a final project completed by the first two authors during Cynthia Dwork’s course “Topics in Theory for Society: The Theory of Algorithmic Fairness” at Harvard (Spring 2024). We thank Cynthia Dwork for initial discussions regarding multicalibration and its applications to complexity and hardness amplification. We thank the anonymous Eurocrypt reviewers for their suggestions and feedback.
@@
The indistinguishability parameter (i.e. “$\gamma$” if two random variables are $\gamma$-indistinguishable) of $X^{k(n)}$ and $Y^{k(n)}$ is upper bounded by the sum of the indistinguishability parameter of $X^{k(n)},\widetilde{X}^{k(n)}$, of $\widetilde{X}^{k(n)},\widetilde{Y}^{k(n)}$, and of $Y^{k(n)},\widetilde{Y}^{k(n)}$. Therefore, $X^{k(n)}$ and $Y^{k(n)}$ are $\left(\frac{1}{4}+k(n)\cdot(n^{-\omega(1)}+\varepsilon(n))\right)$-indistinguishable, which is $\leq 1/2$ as $k(n)=n^{O(1)}$. Thus, $X\equiv^{cS}_{k}Y$.

### v5

@@ 9.1 Acknowledgements
This project was inspired by a final project completed by the first two authors during Cynthia Dwork’s course “Topics in Theory for Society: The Theory of Algorithmic Fairness” at Harvard (Spring 2024). We thank Cynthia Dwork for initial discussions regarding multicalibration and its applications to complexity and hardness amplification. We thank the anonymous Eurocrypt and CCC reviewers for their suggestions and feedback, and we thank Pranay Tankala for bringing to our attention errors in our cited version of the Multicalibration theorem.
@@
The indistinguishability parameter (i.e. $\gamma$” if two random variables are $\gamma$-indistinguishable) of $X^{k(n)}$ and $Y^{k(n)}$ is upper bounded by the sum of the indistinguishability parameter of $X^{k(n)},\widetilde{X}^{k(n)}$, of $\widetilde{X}^{k(n)},\widetilde{Y}^{k(n)}$, and of $Y^{k(n)},\widetilde{Y}^{k(n)}$. Therefore, $X^{k(n)}$ and $Y^{k(n)}$ are $\left(\frac{1}{4}+k(n)\cdot(n^{-\omega(1)}+\varepsilon(n))\right)$-indistinguishable, which is $\leq 1/2$ as $k(n)=n^{O(1)}$. Thus, $X\equiv^{cS}_{k}Y$.